I never get how to do these!! They so frustrating! I believe that there's a formula involved, though I'm not sure...
Each interior angle of a regular polygon measures 156°. How many sides does the polygon have?
A: 13
B: 14
C: 15
D: 16
It's not 14, I can tell you that. I don't know why it's not 14 though.
thank you!! :)

Answer by Alan3354(30993)

I never get how to do these!! They so frustrating! I believe that there's a formula involved, though I'm not sure...
Each interior angle of a regular polygon measures 156°. How many sides does the polygon have?
A: 13
B: 14
C: 15
D: 16
It's not 14, I can tell you that. I don't know why it's not 14 though.
-------------------
The sum of the interior angles of a polygon (all polygons) is
180º*(n-2) (n is the number of sides)
So, each angle is 180*(n-2)/n
----------
156 = 180(n-2)/n
156n = 180(n-2)
156n = 180n - 360
-24n = -360
n = 15
15 sides, 15 angles